# Taylor’s hypothesis and its impact on flux measurements in a forest clearcut

### v1

@@ 3.2.2 DTS temperature spectra
To test whether this procedure generates meaningful results, we used the DTS dataset of temperature, since that encompasses both temporal and spatial information. First, we computed the spectra of temperature fluctuations in a temporal domain, where the frequencies were converted to wavenumbers ($\kappa=2\pi f/U_{C}$) using the bulk convective speed $U_{C}$ (see the blue squares in Fig. 6). Second, the spectra of temperature fluctuations were computed from the spatial data, where the wavenumbers ($\kappa=2\pi/\lambda$, where $\lambda$ is the wavelength) were estimated directly from the spatial distances (red squares in Fig. 6). Third, the temporal spectra of temperature were replotted using the scale-dependent convective speeds to convert the frequencies to wavenumbers (see Eq. 11) while the spectral densities were calculated using the exponential bin framework (black squares in Fig. 6). In Fig. 6, we show these three comparisons for all the six sections of the DTS cable. As noted before, Sections 1-5 describe the along-wind periods (Figs. 6a–e) while Section 6 describes the perpendicular-wind periods (Fig. 6f). Notwithstanding the sections, the spectra in temporal space agree exceptionally well with the spatial spectra, when the frequencies are converted to wavenumbers using Eq. 11. More importantly, a clear $-5/3$ spectral slope is recovered in the temporal domain under this transformation.
3.2.3 EC spectra and cospectra
Figure 7: The ensemble-averaged (a–b) spectra and (c–d) cospectra of scalars (temperature, moisture, and $\rm CO_{2}$) and velocity ($u$ and $w$) components are shown from the clearcut EC measurements, pertaining to the cases when the wind was blowing parallel to the forest edge (along-wind). In (b) and (d), the conversion of frequency to wavenumber is carried out using the scale-dependent convection speeds, as estimated from the DTS data. (e)–(h) The same information as in (a)–(d), but for the perpendicular-wind cases. The spectral and cospectral slopes of $-1$, $-5/3$, and $-7/3$ are shown as solid lines on the respective panels.
After establishing this, we do the same exercise with the EC measurements. Here, we focus separately on the along-wind (Figs. 7a–d) and perpendicular-wind periods (Figs. 7e–h). The spectral and cospectral densities ($S_{xx}$ and $S_{wx}$, where $x=u,w,T,q,c$) in Fig. 7 are normalized with the respective variances and product of the standard deviations. If the mean wind speed from the sonic anemometer ($U$) is used to convert the $f$ values to $\kappa$, the spectra of all the three scalars (temperature, moisture, and $\rm CO_{2}$) and streamwise velocity ($u$) follow a $-1$ spectral slope at lower wavenumbers (Fig. 7a). Regarding $w$, the $-1$ slope is not clearly evident, which is consistent with the attached eddy hypothesis (Townsend, 1976). For the scalars, especially moisture and $\rm CO_{2}$, one could argue that at larger wavenumbers the spectra do not exactly follow the $-5/3$ slope, which could be due to attenuation of the high frequency fluctuations in the gas analyzer sampling tube and instrument noise. However, under the transformation with Eq. 11 ($\kappa z=2\pi fz/U_{C}(\kappa z)$), an extensive region of $-5/3$ slope (nearly two decades, $10.01<\kappa z<$ 1) is recovered for $u$ and the scalars (Fig. 7b).
In fact, Eq. 11 modifies the wavenumber ranges up to which the scaling laws hold. For instance, it shrinks the extent of $-1$ spectral slope while inflating the regions of $-5/3$ (see Fig. 7b). Although in Fig. 7b we show the ensemble-averaged results, if one looks at individual 30-min periods then too this extended $-5/3$ region is conspicuous (see Fig. S5 in the Supplementary Material). For illustration purposes, $u$ and $T$ signals are used as examples in Fig. S5. Upon applying the Kolmogorov’s similarity model on these transformed spectra, one could essentially recompute the dissipation rates of turbulence kinetic energy and scalar variances. This is in accord with previous studies, demonstrating that the dissipation rate estimations from temporal turbulence spectra are sensitive to whether the Taylor’s hypothesis holds (Dahm and Southerland, 1997; Hsieh and Katul, 1997).
Unlike scalars, for the vertical velocity, scale-dependent convective speeds modify the $-5/3$ spectral slope. We explain this anomaly in Appendix C. If we turn our attention towards the flux cospectra (heat, moisture, momentum, and $\rm CO_{2}$), then under the assumption of Taylor’s hypothesis, the cospectral amplitudes at larger wavenumbers show a slope that is nearly equal to $-7/3$ (Fig. 7c). The cospectral slope of $-7/3$ is consistent with the dimensional result of Wyngaard and Coté (1972). For momentum and $\rm CO_{2}$ transport we show their absolute cospectral densities, since their original values are negative. This is evident from the premultiplied cospectra shown in Fig. S6 of the Supplementary Material. Now, if the conversion to wavenumbers is carried out using Eq. 11, we note that the flux cospectral densities display a $-5/3$ scaling instead of $-7/3$ (Fig. 7d). The exact cospectral slope has been debatable since their exponents have shown to be not unique and depend on how the dimensional analysis is performed (Cheng et al., 2020). These same results appear to be true even for the perpendicular-wind cases (Figs. 7e–h). Inspite of such interesting findings, the estimations of the revised $\kappa z$ values and subsequently the spectra are hinged upon how accurate the DTS measurements are. While we discuss this aspect next, we also provide a recipe on how to correct the EC flux measurements based on the results presented so far.
3.3 Implications for EC fluxes
3.3.1 Limitations of DTS
Peltola et al. (2021) showed that at smaller scales of the flow the DTS measurements could be affected by the presence of a white noise. Since white noise has a flat spectrum, in a premultiplied form, it displays a $+1$ spectral slope. To check this, we first computed the DTS spectra of temperature fluctuations from temporal data, and compared the same with EC measurements. To aid the comparison, the spectral amplitudes are normalized by the temperature variances ($\sigma^{2}_{T}$) estimated from the EC measurements and the frequencies are scaled with the standard surface layer scaling ($fz/U$) of Kaimal et al. (1972). Here, we focus only on the along-wind periods although the conclusions remain same for the perpendicular-wind periods too. From Fig. 8a, one could see that the DTS temporal spectra agree exceptionally well with the EC temperature spectra at lower frequencies. However, since the temporal resolution of DTS data ($2.1\ \rm s$) is lower than the EC one ($0.1\ \rm s$), it does not fully capture the fine scale features. Contrary to Peltola et al. (2021), we do not find any evidence of a white noise since there is no presence of a $+1$ spectral slope at larger frequencies. This is likely since our analyses focused on periods with high temperature variance and hence on periods with high signal-to-noise ratio.
Figure 8: For the along-wind cases, (a) the ensemble-averaged pre-multiplied temporal spectra of temperature ($fS_{TT}(f)$) from the DTS data are compared against the EC measurements of temperature. The spectral amplitudes are normalized with the variance of EC temperature ($\sigma^{2}_{T}$) while the frequencies are scaled with the canonical surface layer scaling. (b) The spatial spectra from the DTS data are shown for the along-wind scenario. The $-2/3$ and $+1$ slopes denote the predictions from the Kolmogorov’s spectral model and a pure white-noise case. The vertical dash-dotted line indicates the location $\kappa z=3$. (c)–(d) The impact of using scale-dependent convection speeds instead of a constant one is shown for the heat and moisture cospectra. Here, the scale-dependent convection speeds are estimated from our DTS measurements (pink lines) and from the Cheng et al. (2017) data (blue lines). The red and magenta lines denote when the conversion was carried out using mean wind speed from the EC data ($U$) and bulk convective speed ($U_{C}$) estimated from the space-time correlation curves. (e) For all such different convection speeds, frequencies are plotted against the converted wavenumbers. The horizontal dash-dotted line indicates the critical wavenumber $2\pi/0.3\rm\ m^{-1}$ while the vertical one indicates $f_{\rm cr}$ (Eq. 15). (f) The scatter plot between the surface energy balance components is shown. Here, $\rm R_{n}$ and $\rm G_{0}$ are the net-radiation and soil heat flux components while $\rm SHF$ and $\rm LHF$ represent the sensible and latent heat fluxes in $\rm W\ m^{-2}$. The red squares represent the original values of the sensible and latent heat fluxes and the blue dash-dotted line indicates the best fit. The pink squares represent the values after applying Eq. 17 on the sensible and latent heat fluxes.
On the other hand, if the same information is seen from spatial domain (Fig. 8b), then the DTS premultiplied temperature spectra do show a clear $-2/3$ spectral slope up to wavenumbers $\kappa z\leq 3$ (the vertical dash-dotted line in Fig. 8b). Beyond that, a spurious increase is noted in the spectra although that increase cannot be entirely described as a pure white noise. Nevertheless, this raises a question about the accuracy of $U_{C}(\kappa z)$ values in Fig. 5a. To check that, we carried out a randomly-shuffled exercise on the DTS temperature data. At each instant of time, we randomly shuffled the spatial DTS data, thereby destroying any coherence between them. For randomly-shuffled DTS data, the phase differences between the spatial temperature signals at time instants $t$ and $t+\Delta t$ would be equal to $\pi/2$ at all the wavenumbers since no coherence is expected between them. Therefore, Eq. 5 suggests that for randomly-shuffled signals, $U_{C}(\kappa)$ would vary as ${(\kappa)}^{-1}$, since both $\varphi$ and $\Delta t$ remain constant irrespective of different $\kappa$ values.
This is indeed the case as shown in Fig. S7 of the Supplementary Material. As opposed to the real signals (Fig. S7a), for the randomly-shuffled samples, one can clearly see that the $U_{c}(\kappa z)$ values vary as ${(\kappa z)}^{-1}$ instead of ${(\kappa z)}^{-0.8}$ (Fig. S7b). For clarity purposes, we use a linear $x$ axis in Fig. S7 and restrict the wavenumbers up to $\kappa z\leq 3$. This outcome provides confidence in our estimates of $U_{c}(\kappa z)$ in Fig. 5a. We also note that the exponent $-0.8$ is relatively close to $-1$ and hence our results can be partly affected by the noise seen in Fig. 8b. Nevertheless, in Fig. S8 of the Supplementary Material, we compare three scaling relationships (${(\kappa z)}^{-1/3}$, ${(\kappa z)}^{-0.8}$, and ${(\kappa z)}^{-1}$) for the perpendicular-wind cases, and it is conspicuous that at wavenumbers $\kappa z>1$, $U_{C}(\kappa z)$ values are best described by the relationship ${(\kappa z)}^{-0.8}$.

### v2

@@ 3.2.3 DTS temperature spectra
To test whether this procedure generates meaningful results, we used the DTS dataset of temperature, since that encompasses both temporal and spatial information. For illustration purposes, corresponding to Section-1 of the DTS cable, in Fig. 6, we show the temperature spectra for both the temporal (Fig. 6a) and spatial (Fig. 6b) domains. For the spatial spectra in Fig. 6b, the wavenumbers were computed directly from the DTS spatial separations by using $\kappa=2\pi/\lambda$, where $\lambda$ is the wavelength. In Fig. 6a, the blue lines with square markers indicate the spectra of temperature fluctuations in a temporal domain, where the frequencies were converted to wavenumbers ($\kappa=2\pi f/U_{C}$) using the bulk convective speed $U_{C}$. The temporal spectra in Fig. 6a were replotted with scale-dependent convective speeds, where Eq. 10 was used to convert the frequencies to wavenumbers while the spectral densities were calculated using Eq. 11 (black lines with square markers). Mathematically, this conversion can be expressed as,
$\frac{E_{TT}(2\pi fz/U_{c})}{d(2\pi fz/U_{c})}\equiv\frac{{E_{TT}\left(\frac{\kappa z}{{(1+\kappa z)}^{\alpha}}\right)}\left(1+(1-\alpha)\kappa z\right)}{{{(1+\kappa z)}^{\alpha+1}}d(\kappa z)},$ (20)
where $E_{TT}$ is the spectral energy, computed as the squared magnitudes of the complex Fourier coefficients of temperature. By using $\alpha=0.8$ in Eq. 20 and applying Eq. 18, one recovers a converted temporal spectra with a slope of $5/3+(5/3-1)\times 0.8\approx-1.13$ in the inertial subrange scales (see the magenta line in Fig. 6a). Upon comparison with Fig. 6b, it is clear that the spectra in the spatial domain too show this enhanced slope at larger wavenumbers (see the magenta line in Fig. 6b).
Figure 7: (a)–(f) The converted DTS temporal spectra and spatial spectra are compared for the DTS cable sections, spanning from 1 to 6. Sections 1-5 describe the along-wind periods (a–e) while Section 6 (f) describes the perpendicular-wind periods. The black dash dotted lines in (a)–(f) indicate the 5/3-5/3 spectral slope.
In Figs. 7a–f, we show the comparisons between the converted temporal spectra (gray squares) and spatial spectra (red squares) for all the six sections of the DTS cable. As noted before, Sections 1-5 describe the along-wind periods (Figs. 7a–e) while Section 6 describes the perpendicular-wind periods (Fig. 7f). The error bars denote one standard deviation from the ensemble mean. Notwithstanding the sections, the spectra in temporal space agree well with the spatial spectra, when the frequencies are converted to wavenumbers using Eqs. 10 and 11. While carrying out such conversion, we use $\alpha=0.8$ in Eq. 11. After establishing this, we do the same exercise with the EC measurements. However, instead of restricting ourselves to a specific value of $\alpha$, we carry out the spectral conversion on EC dataset for a range of $\alpha$ values. This is done to systematically investigate the impact of scale-dependent conversion on the EC measurements of turbulence statistics.
3.3 Impact on EC measurements
3.3.1 EC spectra
Figure 8: For the along-wind periods, the ensemble averaged EC spectra of (a) $u$, (b) $w$, (c) $T$, and (d) heat cospectra (wTwT) are shown. The different colored lines indicate the temporal spectra being converted according to a specific exponent of scale-dependent convection speeds ($\alpha$), which vary from 0 to $0.8$ at increments of $0.1$. Here $\alpha=0$ indicates that the conversion has been achieved through the Taylor’s hypothesis. The black lines in (a)–(c) indicate the spectral slope of 5/3-5/3 while the red line indicates a slope of $5/3+0.8\times 2/3\approx-1.13$ (see Eq. 19). For (d), the black line indicates a 7/3-7/3 cospectral slope while the red line indicates a slope of $7/3+0.8\times 4/3\approx-1.3$ (see Eq. 19). For the same along-wind component, in (e)–(h), the flux cospectra of momentum ($uw$), heat ($wT$), moisture ($wq$), and $\rm CO_{2}$ (wcwc) are shown in premultiplied form. Similar to (a)–(d), the different colored lines indicate the range of $\alpha$ values. For all the panels, the spectral and cospectral amplitudes are normalized by respective variances and the product of the standard deviations.
In Figs. 8a–c, we show the ensemble averaged $u$, $w$, and $T$ spectra from the EC dataset, corresponding to the periods when the wind was blowing parallel to the forest edge (along-wind scenario). The conclusions remain the same for the perpendicular-wind periods too, and therefore not being shown. The different colored lines in Figs. 8a–c represent the spectra being converted from temporal to spatial domain using scale-dependent convection speeds (see the legend in Fig. 8a). In particular, we use the following relationship to describe the scale-dependent convection speeds, $U_{C}(\kappa z)/U_{C}=1/(1+\kappa z)^{\alpha}$, with $0\leq\alpha<1$. We choose a range of $\alpha$ values varying between 0 to 0.8 at an increment of 0.1. These conversions are carried out following the Eqs. 9 and 11. Note that $\alpha=0$ indicates the situation with no scale-dependency and hence equivalent to what one would obtain by applying Taylor’s hypothesis. While conducting this analysis, we do consider that the scale-dependent convection speeds have a power-law relationship at smaller scales of the flow. Accordingly, these spectral conversions are sensitive to $\alpha$ values rather than whether $U$, $U_{C}$, or $U_{e}$ were used to scale $U_{C}(\kappa z)$.
At a first glance, these spectral results in Figs. 8a–c appear counter-intuitive. This is because, the temporal spectra, without any application of scale-dependent conversion, follow a clear 5/3-5/3 slope (see the black dash-dotted lines in Figs. 8a–c). In accordance with Eq. 19, any non-zero $\alpha$ values would yield a spectral slope of $\beta+(\beta-1)\alpha$ in the inertial subrange scales, where $\beta$ (=5/3=5/3 for our case) is the slope in the frequency domain ($\alpha=0$). Since $5/3>1$, the term $(\beta-1)\alpha>0$, and hence the spectral slopes will exceed 5/3-5/3 when the scale-dependent conversion is carried out, irrespective of whether $\alpha$ is $1/3$ or $0.8$. This is precisely what we observe in Figs. 8a–c that the spectra become progressively flattened as $\alpha$ increases. In that respect, we agree with cheng2017failure that applying Taylor’s hypothesis systematically underestimates the turbulent spectrum in the inertial subrange. Specific to $\alpha=0.8$, which is the value we obtain for our dataset (see Fig. 5b), the converted spectral slopes are equal to $5/3+0.8\times 2/3\approx-1.13$ (see the red dash-dotted lines in Figs. 8a–c).
It is prudent to ask, what does it physically mean for the converted temporal spectra, at an $\alpha$ value of 0.8, to show a broad region of $\approx-1.13$ spectral slope with no 5/3-5/3 region?. As shown in Figs. 6b and 7, the spatial spectra of DTS temperature fluctuations show a similar enhanced slope ($\approx-1.13$) at the inertial subrange wavenumbers ($\kappa z\geq 1$). Since the DTS dataset is limited by temperature information only, we do not have equivalent spatial information for the velocity counterpart. In the present scenario, by assuming that the turbulence structures in the velocity fields follow the same scale-dependent convection speed relationship as of temperature, we consider that the same spectral enhancement at the inertial subrange scales occur for $u$ and $w$ components. This of course needs to be validated through other means than DTS measurements (perhaps large eddy simulations) but that is beyond the scope of this work. Keeping such limitation in mind, we explore two potential reasons for this spectral enhancement.
3.3.2 Dimensional analysis
The first reason is small-scale intermittency that could cause deviations from the expected 5/3-5/3 spectral slope (dubrulle1994intermittency). However, such corrections are small for the second-order turbulence statistics (i.e. spectra) and therefore cannot explain our observations. The second reason is turbulence anisotropy imparted by the large-scale motions to smaller scales of the flow (chowdhuri2024quantifying). To account for such large‐scale inertial scale interactions, katul2025scaling proposed a dimensional analysis framework by considering an additional influence of large-scale field on the inertial subrange turbulence statistics (see their Eq. 15). Although katul2025scaling developed this framework for pressure fluctuations in near-neutral roughness sublayer flows, we extend the same to our case. To start with, let us consider that the turbulence velocity spectra in the inertial subrange scales depend on the large-scale velocity field, which we denote as $X$. For the scalar field, such as temperature, given the strong relationship between the sweeping speed of temperature structures and TKE of the flow (see Fig. 4e), we extend the same dependence on $X$ to the temperature spectra. Therefore, the spectral amplitudes of velocity and temperature in the inertial subrange scales have a functional dependence of the form,
$\begin{split}S_{xx}(\kappa)=F(\epsilon_{\rm TKE},\kappa,X)\\ S_{TT}(\kappa)=F(\epsilon_{\theta},\epsilon_{\rm TKE},\kappa,X).\end{split}$ (21)
In Eq. 21, $x=u,w$, $\epsilon_{\theta}$ and $\epsilon_{\rm TKE}$ are the dissipation rates of temperature variance and turbulence kinetic energy and $\kappa$ is the wavenumber. We now conduct a dimensional analysis and for that purpose, we consider the three fundamental dimensions $[L,T,K]$, where $L$, $T$, and $K$ represent the dimensions of length, time, and temperature, respectively. Since the dimensions of $S_{xx}(\kappa)$ and $S_{TT}(\kappa)$ are $L^{3}T^{-2}$ and $K^{2}L$, respectively, one could write Eq. 21 in its dimensional form as,
$\begin{split}L^{3}T^{-2}\sim\epsilon^{a_{1}}_{\rm TKE}\kappa^{b_{1}}X^{c_{1}}\\ K^{2}L\sim\epsilon^{a_{2}}_{\theta}\epsilon^{b_{2}}_{\rm TKE}\kappa^{c_{2}}X^{d_{2}},\end{split}$ (22)
where the dimensions of $\epsilon_{\theta}$, $\epsilon_{\rm TKE}$, $\kappa$, and $X$ are, $K^{2}T^{-1}$, $L^{2}T^{-3}$, $L^{-1}$, and $LT^{-1}$, respectively. Henceforth, we remove the subscript TKE from $\epsilon_{\rm TKE}$. By substituting these dimensions in Eq. 22, we get,
$\begin{split}L^{3}T^{-2}\sim L^{2a_{1}-b_{1}+c_{1}}T^{-(3a_{1}+c_{1})}\\ K^{2}L\sim K^{2a_{2}}L^{2b_{2}-c_{2}+d_{2}}T^{-(a_{2}+3b_{2}+d_{2})},\end{split}$ (23)
By equating the similar dimensions at both sides, and after some algebraic manipulations, one ends up with the following spectral forms,
$\begin{split}S_{xx}(\kappa)\sim{\epsilon}^{-(1+c_{1})}\kappa^{c_{1}}X^{5+3c_{1}}\\ S_{TT}(\kappa)\sim\epsilon_{\theta}{\epsilon}^{-(2+c_{2})}\kappa^{c_{2}}X^{5+3c_{2}}.\end{split}$ (24)
Now, in the place of $X$, we put the velocity scale ${(\rm TKE)}^{1/2}$, which represents the energetic large-scale motions, and further rewrite Eq. 24 as,
$\begin{split}S_{xx}(\kappa)\sim{\epsilon}^{-(1+c_{1})}\kappa^{c_{1}}{\rm TKE}^{(5+3c_{1})/2}\\ S_{TT}(\kappa)\sim\epsilon_{\theta}{\epsilon}^{-(2+c_{2})}\kappa^{c_{2}}{\rm TKE}^{(5+3c_{2})/2}.\end{split}$ (25)
If $c_{1}=c_{2}=-5/3$ in Eq. 25, one recovers the standard Kolmogorov spectra of temperature and velocity,
$\begin{split}S_{xx}(\kappa)\sim{\epsilon}^{2/3}\kappa^{-5/3}\\ S_{TT}(\kappa)\sim\epsilon_{\theta}{\epsilon}^{-1/3}\kappa^{-5/3},\end{split}$ (26)
with no dependence whatsoever on the velocity scale ${\rm TKE}^{1/2}$. However, if the spectral slopes are of the form $c_{1}=-5/3+\delta_{1}$ and $c_{2}=-5/3+\delta_{2}$, then Eq. 25 dictates,
$\begin{split}S_{xx}(\kappa)\sim{\epsilon}^{2/3-\delta_{1}}\kappa^{-5/3+\delta_{1}}{(\rm TKE)}^{3\delta_{1}/2}\\ S_{TT}(\kappa)\sim\epsilon_{\theta}{\epsilon}^{-(\delta_{2}+1/3)}\kappa^{-5/3+\delta_{2}}{(\rm TKE)}^{3\delta_{2}/2}.\end{split}$ (27)
The non-vanishing terms ${(\rm TKE)}^{3\delta_{1}/2}$ and ${(\rm TKE)}^{3\delta_{2}/2}$ in Eq. 27 imply that the presence of a large-scale velocity field can indeed cause deviations from the standard Kolmogorov spectral form. However, unlike the Kolmogorov spectra, in this case, the exponents $c_{1}$ and $c_{2}$ are not unique (hence not universal) and $c_{2}\approx-1.13$ emerges to be one of the possibilities. Apart from that, in order to satisfy Eq. 25, the exponents $c_{1}$ and $c_{2}$ need not be equal for the velocity and temperature signals. Therefore, by introducing the effects of the large-scale velocity field on inertial subrange scale motions positions us to have spectral exponents larger than 5/3-5/3 with differences being allowed to exist between the velocity and temperature signals. We clarify that our intent behind conducting this dimensional analysis is to provide a plausible reason for enhanced spectral slopes at inertial subrange scales, rather than deriving a new spectral model of velocity and temperature from first principles. For instance, it is currently unclear up to what wavenumber limit such enhancements in spectral slopes continue. Moreover, are the values of $c_{1}$ and $c_{2}$, satisfying Eq. 25, equal for velocity and temperature signals? To answer these questions more spatially explicit field measurements and numerical simulations are required, which is a matter of future research.
3.3.3 EC cospectra
Next, we investigate the impact of $\alpha$ on flux cospectra. In Fig. 8d we show the heat flux ($wT$) cospectra, normalized with the product of the standard deviations of vertical velocity ($\sigma_{w}$) and temperature ($\sigma_{T}$). Analogous to Figs. 8a–c, the heat flux cospectra have been transformed using Eqs. 9 and 11 with different $\alpha$ values. The $wT$ cospectral amplitudes, without any application of scale-dependent conversion, closely follow a 7/3-7/3 slope at larger wavenumbers (see the black dash-dotted line in Fig. 8d). As Eq. 19 predicts, this slope transforms to $7/3+(7/3-1)\times 0.8$ for $\alpha=0.8$. The exact cospectral slope has been debatable since their exponents have shown to be not unique and depend on how the dimensional analysis is performed (cheng2020power). Therefore, in that spirit, it is possible to incur a cospectral slope of $7/3+(7/3-1)\times 0.8$ in the spatial domain, which is nearly equal to 1.3-1.3 (see the red dash-dotted line in Fig. 8d). We extend this analysis to other fluxes, namely momentum, moisture and $\rm CO_{2}$.
It is natural to plot the cospectra in a premultiplied form since the area under this log-linear coordinate system is directly related to the flux contributions across different scales (kaimal1994atmospheric). In Figs. 8e–h, we show the premultiplied cospectra of momentum ($uw$), heat ($wT$), moisture ($wq$), and $\rm CO_{2}$ (wcwc) vertical fluxes. Similar to Figs. 8a–d, the different colored lines indicate different $\alpha$ values. As one would expect, the slopes of the premultiplied cospectra increase as $\alpha$ increases. Quite interestingly, the behavior of $uw$ (Fig. 8e) and $wc$ (Fig. 8h) cospectra remain different than heat (Fig. 8f) and moisture (Fig. 8g). It appears that the rate by which the cospectral slopes increase with $\alpha$ values remain smaller for momentum and $\rm CO_{2}$ fluxes. This occurs because, these results are ensemble averaged and the run-to-run variations in $uw$ and $wc$ cospectra are larger than what are present in the heat and moisture cospectra (not shown). Since we apply our spectral conversion technique on individual 30-min periods, the run-to-run variations could have impacted the results when they are ensemble averaged.
Notwithstanding such differences, it is clear that the scale-dependent conversion increases the slopes of cospectral curves. Particularly, the results from Figs. 8f and g demonstrate that non-zero $\alpha$ values can cause an appreciable increase in the heat and moisture fluxes at smaller scales of the flow. As conjectured by cheng2017failure, such enhancement could in fact increase the flux values when Eq. 19 is integrated across the entire inertial subrange scales and by doing so perhaps one could solve the enigma of surface energy imbalance (mauder2020surface). This remains a topic for our future research. However, as we show below, the frequency to wavenumber conversion through Eq. 9 yields a critical frequency limit beyond which the space-time conversion is uncertain.